A large weather balloon whose mass is 222 kg is filled with helium gas until its volume is 322 m3. Assume the density of air is 1.20 kg/m3 and the density of helium is 0.179 kg/m3.

(a) Calculate the buoyant force acting on the balloon.

          3786.72 N

(b) Find the net force on the balloon.

                      N upward

(c) What additional mass can the balloon support in equilibrium?

                     kg

Please help!
Answer #1

here mb = 222 kg V = 322 m3   pa = 1.20 Kg / m3 ph = 0.179 Kg / m3

as we know that the buoyant force

Fb = pagV = 3786.72 N

(b) net force on the balloon

Fnet = (pa - ph)gV - mg

= (1.20 - 0.179) x 9.81 x 322 - 222 x 9.81

= 1047.335 N Ans

(c) for the equilibrium condition

Fbuoyant = Fgravitational

3786.72 = mb g

mb = 3786.72 / 9.81

= 386.006 Kg Ans